OpenStudy (amonoconnor):

Find the Differential of: "y = ln[sqrt.(1+x^2)]" Any and all help is greatly appreciated!

zepdrix (zepdrix):

Hmm so where you stuck? :d

zepdrix (zepdrix):

\[\large\rm y=\ln\left[\sqrt{1+x^2}\right]\]

zepdrix (zepdrix):

Hmm I don't see a t anywhere :d

OpenStudy (amonoconnor):

Lols, sorry. My book has t's. I was doing x's on here to make it "easier". My bad. yeah, that's supposed to be an "x" ;)

zepdrix (zepdrix):

broski. I strongly recommend you commit this derivative to memory. It shows up sooo often, its really worth a spot in your brain.\[\large\rm \frac{d}{dx}\sqrt{x}=\frac{1}{2\sqrt x}\] The derivative of square root is `one over two square roots`. Don't bother with power rule on that guy.

zepdrix (zepdrix):

\[\large\rm y=\ln\left[\sqrt{1+x^2}\right]\]First step, derivative of log, and "set up" chain,\[\large\rm dy=\frac{1}{\sqrt{1+x^2}}\cdot d\sqrt{1+x^2}\]

OpenStudy (amonoconnor):

Gotcha! I did that:)

zepdrix (zepdrix):

Apply square root derivative, and "set up" the next chain,\[\large\rm dy=\frac{1}{\sqrt{1+x^2}}\cdot \frac{1}{2\sqrt{1+x^2}}\cdot d (1+x^2)\]

OpenStudy (amonoconnor):

Wait... the last term there. The d(1+x^2).. what's that? Is that dx?

zepdrix (zepdrix):

Hmm no. We're not to dx just yet :)

OpenStudy (anonymous):

you can simplify the expression.. then it would be quite easy ln x^a = a ln x that will remove the sq rt and differentiation will become easy

zepdrix (zepdrix):

Yes, that's also a good idea :) Saves you a couple steps.

zepdrix (zepdrix):

As second step, you took this derivative:\[\large\rm d\sqrt{stuff}\quad=\quad \frac{1}{2\sqrt{stuff}}\cdot d(stuff)\]Gotta differentiate that stuff on the inside by chain rule, ya?

OpenStudy (amonoconnor):

I'm good on the first two pieces of that... but why do we need to take the derivative of (sqrt.(1+x^2) twice?

zepdrix (zepdrix):

We took the derivative of log first. Then square root second. Then (1+x^2) third.

OpenStudy (amonoconnor):

so... \[d (\ln(stuff)) = 1/(stuff) * Chain \Rule of net stuff * d(stuff)\] ?

OpenStudy (amonoconnor):

Gotcha... Okay..

zepdrix (zepdrix):

And then I guess derivative of the inner most function gives us,\[\large\rm dy=\frac{1}{\sqrt{1+x^2}}\cdot \frac{1}{2\sqrt{1+x^2}}\cdot (2x)dx\]Ya? :o

OpenStudy (amonoconnor):

Yeppers, I'm there:)

zepdrix (zepdrix):

And then uhhh, simplify if you can.

OpenStudy (amonoconnor):

I think I see now... I didn't know about doing the chain rule last half part err... "twice". Damn, my prof is so.... irritating. He never does examples like we have in the homework!! >:/

zepdrix (zepdrix):

Oh you have one of those? :C grr

OpenStudy (amonoconnor):

That makes sense though, I get it now:)

OpenStudy (amonoconnor):

Where do you go to school? :)

zepdrix (zepdrix):

Divu's tip was quite helpful if you want to try that one at some point.\[\large\rm \ln\left[(1+x^2)^{1/2}\right]\quad=\quad \frac{1}{2}\ln\left[1+x^2\right]\]And that will save you the work of one of the chains.

zepdrix (zepdrix):

University of Central Florida. I hate it :o school is so harrrrd.

OpenStudy (amonoconnor):

True, that looks slick. Oh, nice! My friend goes to Emery-Riddle School of Aviation down in Florida, and I've been wishing I could be in that weather right now. I go to the University of Minnesota, Twin Cities, and it's been like, 50-60 degrees... I'm missing my shorts and tan already :(

zepdrix (zepdrix):

Emery? :o Ooo nice school. Yah the weather can get pretty rough down here sometimes... I had to throw an extra shirt on due to a light breeze. Brrrr! I think it was in the low 70s :(

OpenStudy (amonoconnor):

Haha, low 70's;) If only... I wish it would be that nice even through November! I've been thinking about snow, and where I'm going to put my bike where it won't get drifted in between classes! Yeah, Emery Riddle sounds super sweet. My friend was flying a plane on literally, the first day of class!! It's insane, and also pricey..

zepdrix (zepdrix):

Woahhh :O fun!

OpenStudy (amonoconnor):

Alright, so if I have: \[y = \frac{1-v^2}{1+v^2}\] And I'm looking to get the "differential" again, is quotient rule the best bet?

zepdrix (zepdrix):

Umm.. quotient rule is fine I guess. I would prefer long division though.\[\rm \frac{1-v^2}{1+v^2}\quad=\quad \frac{-(v^2-1)}{v^2+1}\quad=\quad \frac{-(v^2+1-2)}{v^2+1}\quad=\quad -\frac{v^2+1}{v^2+1}+\frac{2}{v^2+1}\] And then from this point,\[\large\rm =-1+2(v^2+1)^{-1}\]You can simply apply power rule. If those Algebra steps are way confusing though, you can just stick with quotient ^^

zepdrix (zepdrix):

That little trick works way better on problems like this:\[\large\rm \frac{x^2}{x^2+1}\]Maybe it's a little confusing for your particular problem, since we had to mess with the negatives so much.

OpenStudy (amonoconnor):

Yeah, that was a lil' over my head.. I'm going to finish doing Quotient Rule. What do you get as the final answer?

zepdrix (zepdrix):

\[\large\rm dy=\frac{-4v^3}{(1+v^2)^2}dv\]

zepdrix (zepdrix):

Sec, checking my work :D

OpenStudy (amonoconnor):

Don't the 2v^3 's cancel out, and there are two (-2v)'s? :/

zepdrix (zepdrix):

Mmm let's just find out I guess :d

OpenStudy (amonoconnor):

:D

OpenStudy (amonoconnor):

I'm like, 90% sure lols.

zepdrix (zepdrix):

\[\large\rm y=\frac{1-v^2}{1+v^2}\] \[\large\rm dy=\frac{(-2v)(1+v^2)-(2v)(1-v^2)}{(1+v^2)^2}dv\]This step ok?

OpenStudy (amonoconnor):

Good there:) I got that too:)

zepdrix (zepdrix):

Ah yes :) I pulled out too many negatives I think. Good call.

zepdrix (zepdrix):

-4v on top then?

OpenStudy (amonoconnor):

Yeppers man:)

zepdrix (zepdrix):

smoooooth! ୧ʕ•̀ᴥ•́ʔ୨

OpenStudy (amonoconnor):

If... \[y = e^{x/10}\] Then... \[dy = \frac{x}{10}lne \] ?

zepdrix (zepdrix):

Hmm where did this magical log come from? I only see one on the right side.

zepdrix (zepdrix):

Remember derivative for e^x? Gives you the same thing back, ya?\[\large\rm \frac{d}{dx}e^{x}=e^{x}\]And we'll generalize it a little more for chain rule,\[\large\rm \frac{d}{dx}e^{stuff}=e^{stuff}~\frac{d}{dx}stuff\]

OpenStudy (amonoconnor):

I remember that, about d/dx of e^x! Ah.. so I don't need to take the natural log of both sides or anything??

zepdrix (zepdrix):

Logarithmic differentiation? Only if `both the base and exponent are x stuff`. \(\large\rm y=f(x)^{g(x)}\qquad \text{<--- Apply Logarithmic Differentiation}\) Example: \(\large\rm y=x^{cos(x)}\)

zepdrix (zepdrix):

I mean, yes you should go through the log differentiation process at least one time to justify the exponential derivative, but after that first time, it's a nice shortcut to remember. Here is the more general rule though:\[\large\rm \frac{d}{dx}a^x=a^{x}\ln(a)\]So NORMALLY, you have to multiply by the log of the base. But ln(e) is just 1, so we generally skip that on the e^x derivative.

OpenStudy (amonoconnor):

Gotcha... so, it should be like this? \[dy = e^{\frac{x}{10}}*\frac{(1*10) - (x*0)}{100}\]

zepdrix (zepdrix):

Yes, you COULD do quotient rule. Lemme state this as I did for the exponential rule: Apply quotient rule only if `the numerator and denominator are both x stuff`.

zepdrix (zepdrix):

Instead do this:\[\large\rm \frac{x}{10}=\frac{1}{10}x\]And apply your power rule or whatever.

zepdrix (zepdrix):

\[\large\rm \frac{d}{dx}ax=a\]ya? :)

OpenStudy (amonoconnor):

True! :)

zepdrix (zepdrix):

\[\large\rm y=e^{x/10}\]\[\large\rm dy=e^{x/10}\cdot d~\left(\frac{x}{10}\right)\]So what do you get for that chain? And final answer?

zepdrix (zepdrix):

Confused? :o

OpenStudy (amonoconnor):

Final answer: .1(e^(x/10))*dx :)

zepdrix (zepdrix):

hmmm

OpenStudy (amonoconnor):

My book says it's right... ? :/

zepdrix (zepdrix):

Oh you're being sneaky, .1 = 1/10, Ok fine fine

OpenStudy (amonoconnor):

yeah, sorry:)

OpenStudy (anonymous):

Interesting thread, can you find the differential for the following?You may have to use a substitution! \[y=\tan^{-1}(\frac{a}{\sqrt{x^2-a^2}})\] \[y=\sin^{-1}(2x \sqrt{1-x^2})\]

OpenStudy (amonoconnor):

I'm sorry, but I am jus trying to get something on paper for my assignment in Calc that's due today... or tomorrow, depends on wether "Today" starts once I sleep, or now, since it's 2am lols. I would love to challenge myself, but I have been stuck on a pellet-ton of problems, and am just trying to get something down for all of them, so I can get some credit.. Sorry:/

OpenStudy (amonoconnor):

Sh1t = pellet... ? why?

zepdrix (zepdrix):

ya the filter word is hilarious lol

OpenStudy (amonoconnor):

1 = i

OpenStudy (anonymous):

Oh best of luck, feel free to try them later if you want to, I think these problems would be in your textbook anyway if you read on further into your book

OpenStudy (amonoconnor):

Thank you!! I certainly will take a peek, once the fire's put out here lols. :)

OpenStudy (shamil98):

zepdrix you are gr8 thanks so much for helping

imqwerty (imqwerty):

@amonoconnor shіt

OpenStudy (amonoconnor):

What's up now @imqwerty? Sorry, I went to bed last night.

imqwerty (imqwerty):

haha i jst tacked the filter :D